Compute the length of the other two legs of a right triangle with a hypotenuse of 5 feet and one angle of 30 degrees
Answer by dnanos(83)

a=5 B=30
We know that if B=30 then the leg b=a/2 in a right triangle.(A=90 degrees).
So b=5/2=2.5 feet
c=sqrt(a^2-b^2) from the Pythagorean theorem.
c=sqrt[5^2-2.5^2)
c=sqrt(25-6.25)
c=sqrt18.75
c~4.33
c~4.33 feet